### v1

# Learning to chain-of-thought with Jensen’s evidence lower bound

Abstract
We propose a way to optimize chain-of-thought with reinforcement learning, but without external reward function. Our algorithm relies on viewing chain-of-thought as latent variable as part of a probabilistic inference problem. Contrary to the full evidence lower bound, we propose to apply a much simpler Jensen’s lower bound, which derives tractable objectives with simple algorithmic components (e.g., without the need for parametric approximate posterior), making it more conducive to modern large-scale training. The lower bound approach naturally interpolates other methods such as supervised fine-tuning and online reinforcement learning, whose practical trade-offs we will illustrate. Finally, we show that on mathematical reasoning problems, optimizing with Jensen’s lower bound is as effective as policy gradient with external reward. Taken together, our results showcase as a proof of concept to this new algorithmic paradigm’s potential to more generic applications.
Machine Learning, ICML
$o$$z$$\theta$$\phi$
(a) Probabilistic inference
$o$$z$$\theta$$\phi$$a^{*}$$c$$\theta$$\phi$
(b) CoT with full ELBO
$o$$z$$\theta$$\phi$$a^{*}$$c$$\theta$$\phi$$a^{*}$$c$$\theta$$\theta$
(c) CoT with Jensen’s bound
$o$$z$$\theta$$\phi$$a^{*}$$c$$\theta$$\phi$$a^{*}$$c$$\theta$$\theta$$a^{*}$$c$$\theta_{\text{ref}}$$\theta$
(d) CoT with Jensen’s bound with KL regularization
Figure 1: Graphical models for various algorithmic formulations discussed in this work. Solid lines represent generative models and dashed lines represent inference models. Circles represent random variables and squares represent parameters. Shading indicates that the random variable is observed, and is used for providing feedback for the learning process. For CoT optimization, $a^{*}$ is a simplified notation for the binary optimality variable $\mathds{1}_{\{a=a^{\ast}\}}$ from the random variable $a$. See Appendix A for a more detailed explanation.
1 Introduction
Chain-of-thought is playing an increasingly important role for large language models. It has been observed that chain-of-thoughts can elicit significant performance improvements from language models, which scales positively with model sizes (Wei et al., 2022). Chain-of-thought can be instantiated in different ways, in the case of reasoning, it usually refers to the step-by-step solution leading up to a final conclusion (Ling et al., 2017; Cobbe et al., 2021; Lightman et al., 2023). Learning to reason with chain-of-thought has proved highly effective, and can endow models with capabilities well beyond pre-training and regular fine-tuning (Jaech et al., 2024; Guo et al., 2025; Lambert et al., 2024; Team et al., 2025).
There are many algorithmic approaches to optimizing chain-of-thoughts for language models, such as prompting, supervised fine-tuning (SFT) and reinforcement learning (RL) (Ling et al., 2017; Cobbe et al., 2021; Wei et al., 2022; Jaech et al., 2024). A rather intriguing perspective is to view chain-of-thought optimization as a latent variable modeling problem. In short, given an initial problem, chain-of-thoughts are interpreted as intermediary latent variables such that the desirable conclusion becomes more likely (Hoffman et al., 2024; Hu et al., 2024). Such an interpretation naturally fits the intuition we have for chain-of-thoughts, that they should serve as intermediary outputs that improve the final prediction. This formulation also brings about the algorithmic developments from the latent variable modeling literature.
In this work, we take a pragmatic approach to adapting methodologies from the latent variable modeling literature. We propose an algorithm based on a single-sample or multiple-samples version of the Jensen’s evidence lower bound, named after the fact that our objective is obtained by simply applying a Jensen’s inequality to the evidence (Jensen, 1906). Optimizing with such a simple lower bound forgoes the complication of variational posterior, which is inherent to the full evidence lower bound (Blei et al., 2017; Kingma and Welling, 2013b; Rezende et al., 2014). This makes our algorithm more suitable for contemporary large-scale training for language model applications, where training auxiliary models is as expensive as the main model (Brown et al., 2020; Achiam et al., 2023).
The final algorithm takes a rather simple form - in a nutshell, it consists of an hybrid online RL-like and SFT-like loss. When viewed as an alternative to online RL, the lower bound method does not require any external reward, unlike regular RLHF algorithms (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). In more details, our technical contributions are as follows:
  • We derive the Jensen’s evidence lower bound in Section 3, the stochastic optimization algorithm, as well as its multi-sample extension (Burda et al., 2015) which tightens the bound in Section 4.
  • We draw insightful connections between the full ELBO, RL with specialized reward and the proposed lower bound algorithm in Section 5. We also show how regularized RLHF as a whole can be coherently interpreted as a latent variable modeling problem. See Figure 1 for a graphical illustration.
  • Besides theoretical equivalence, we highlight the practical trade-offs of applying RL vs. lower bound optimization. The key trade-off factors include the choice of rewards, which we detail in Section 6.
  • Finally in Section 8, we show that optimizing with Jensen’s evidence lower bound and its multi-sample variants is competitive compared to online RL, even without explicit access to a more sophisticated reward function.
2 Reinforcement learning for language model
A language model can be understood as a policy $\pi_{\theta}$ in the context of reinforcement learning. Given a prompt $x$, the policy generates a response $y$, which then gets assessed by a human user. Usually, the objective is to optimize $\pi_{\theta}$ such that certain reward function $r(x,y)$ that captures human preference is maximized (Christiano et al., 2017; Ouyang et al., 2022). Formally, consider the maximization problem
$\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r(x,y)\right]-\beta\mathbb{KL}\left(\pi_{\theta}(\cdot|x),\pi_{\text{ref}}(\cdot|x)\right)$ (1)
with a KL regularization that encourages $\pi_{\theta}$ to stay close to the reference policy. The reward $r(x,y)$ captures the human preference of response $y$ in response to prompt $x$ and can take various forms: for example, it can be extracted from human annotations (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022), generated atuomatic feedback such as code execution (Gehring et al., 2024). We focus on a specialized setting where the reward is derived from access to certain ground truth of the problem.
2.1 RL from ground truth feedback
We focus on applications where the prompt $x$ typically specifies a question and there is an example of an desirable ground truth $a^{\ast}$. Such a formulation includes formally verifiable applications such as mathematical reasoning (Hendrycks et al., 2021; Uesato et al., 2022; Lightman et al., 2023) where $x$ is a question and $a^{\ast}$ is the ground truth answer. Usually, the reward $r(x,y)$ is calculated by matching $a^{\ast}$ against the answer $a$, e.g., using string match.
2.2 Chain-of-thought
For aforementioned applications where the model is required to reason about the question $x$ and generate an answer $a$, getting the model to generate chain-of-thoughts - a sequence of reasoning steps $c$ leading up to the final conclusion (Ling et al., 2017; Wei et al., 2022). Henceforth, we can decompose the generation $y=(c,a)$ into a chain-of-thought $c$ and an answer $a$. The generative process for the response $y\sim\pi_{\theta}(\cdot|x)$ is made more concrete as
$c\sim\pi_{\theta}(\cdot|x),a\sim\pi_{\theta}(\cdot|x,c).$ (2)
Given a prompt $x$, the intuitive role of chain-of-thought is such that it makes the marginal likelihood of the ground truth answer $a^{\ast}$ higher. As such, we can interpret chain-of-thought as a latent variable and formulate the optimization of chain-of-thought as latent variable modeling (Hu et al., 2024; Hoffman et al., 2024).
3 A Jensen’s lower bound for chain-of-thought as latent variable modeling
We start with the initial motivation to increase the marginal likelihood of the ground truth answer $a^{\ast}$ (i.e., the evidence) given the generative process in Eqn (2)
$\max_{\theta}\log\pi_{\theta}(a^{\ast}|x).$ (3)
Directly optimizing the log likelihood is not tractable because its gradient cannot be estimated via samples in an unbiased way (see, e.g., discussion on this in the probabilistic inference literature (Blei et al., 2017)). As the main contribution of this work, we propose a tractable lower bound objective by directly applying the Jensen inequality to lower bound the log likelihood
$\log\pi_{\theta}(a^{\ast}|x)$ $=\log\mathbb{E}_{c\sim\pi_{\theta}(\cdot|x)}\left[\pi_{\theta}(a^{\ast}|x,c)\right]$
$\geq\underbrace{\mathbb{E}_{c\sim\pi_{\theta}(\cdot|x)}\left[\log\pi_{\theta}(a^{\ast}|x,c)\right]}_{\mathcal{L}_{\theta}(x,a^{\ast})},$ (4)
where we exchange the order of the concave $\log$ function and expectation $\mathbb{E}\left[\cdot\right]$. There are conditions under which the lower bound $\mathcal{L}_{\theta}(x,a^{\ast})$ is tight. For example, if all chain of thoughts $c$ in the support of $\pi_{\theta}(\cdot|x)$ induce the same probability of predicting the ground truth answer $\pi_{\theta}(a^{\ast}|x,c)$, i.e., $\pi_{\theta}(a^{\ast}|x,c)=\pi_{\theta}(a^{\ast}|x,c^{\prime}),\forall c,c^{\prime}\in\text{supp}\left(\pi_{\theta}(\cdot|x)\right)$. In practice when the optimization is approximate, such conditions are not likely to hold. As a result, there might be a gap between the lower bound and $\log\pi_{\theta}(a^{\ast}|x)$ and we will examine its empirical impact in practice.
The gap between the marginal log likelihood and the lower bound can be expressed as the KL divergence between $\pi_{\theta}$ and the posterior distribution (Blei et al., 2017)
$\log\pi_{\theta}(a^{\ast}|x)-\mathcal{L}_{\theta}(x,a^{\ast})=\mathbb{KL}\left(\pi_{\theta}(\cdot|x),p^{\pi_{\theta}}(\cdot|x,a^{\ast})\right),$
where $p^{\pi_{\theta}}(c|x,a^{\ast})\coloneqq\frac{\pi_{\theta}(a^{\ast}|x,c)\pi_{\theta}(c|x)}{\sum_{c^{\prime}}\pi_{\theta}(a^{\ast}|x,c^{\prime})\pi_{\theta}(c^{\prime}|x)}$. The posterior defines a distribution over chain-of-thought, and effectively denotes how likely is the chain-of-thought $c$ given that the ground truth answer is $a=a^{\ast}$ and the prompt is $x$. For experienced readers, this lower bound is closely related to the evidence lower bound (Kingma and Welling, 2013a; Blei et al., 2017), which we will elaborate more below.
3.1 Stochastic gradient estimate
The lower bound permits stochastic gradient estimates. Concretely, given samples from the current policy $c\sim\pi_{\theta}(\cdot|x)$, we can construct an estimate of $\nabla_{\theta}\mathcal{L}_{\theta}(x,a^{\ast})$ as
$\underbrace{\log\pi_{\theta}(a^{\ast}|x,c)\nabla_{\theta}\log\pi_{\theta}(c|x)}_{g_{1}}+\underbrace{\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x,c)}_{g_{2}}.$ (5)
The gradient has two terms: $g_{1}$ is a REINFORCE gradient estimate with $\log\pi_{\theta}(a^{\ast}|x,c)$ as the reward function for sampled chain-of-thought $c$ (Thompson, 1933). The second gradient $g_{2}$ is reminiscent of a supervised learning loss that encourages the model to predict ground truth answer $a^{\ast}$ given sampled chain-of-thought $c$.
In practice, we can add a control variate to the REINFORCE gradient estimate to reduce variance. One option is to learn a prompt-answer dependent function (Schulman et al., 2017); another sample-based alternative is to generate $n$ i.i.d. chain-of-thoughts in parallel $c_{i}\sim\pi_{\theta}(\cdot|x)$, and construct leave-one-out control variates $v_{i}=\frac{1}{n-1}\sum_{j\neq i}\log\pi_{\theta}(a^{\ast}|x,c_{j})$ (Mnih and Rezende, 2016; Kool et al., 2019). The overall gradient estimate is the average over $n$ samples:
$\frac{1}{n}\sum_{i=1}^{n}\Big{[}\left(\log\pi_{\theta}(a^{\ast}|x,c_{i})-v_{i}\right)\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$
$+\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x,c_{i})\Big{]}.$
Note the control variates $v_{i}$s do not introduce any bias to the gradient estimate since they are statistically independent from $\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$.
Connections to supervised fine-tuning.
In the very special case where there is no chain-of-thought, the gradient estimate reduces to just the SFT part $\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x)$ which is effectively the supervised fine-tuning loss from prompt $x$ to answer $a^{\ast}$. Here, the key difference is that the loss $\pi_{\theta}(a^{*}|x,c_{i})$ further conditions on the chain-of-thoughts $c_{i}$s whose distribution changes over time and introduce more non-stationarity to the optimization process.
4 Tightening the objective via multi-sample Jensen’s lower bound
If the Jensen’s lower bound is loose, it will induce a sizable discrepancy from the true objective of interest. We need strategies to tighten the lower bound for policy optimization.
A similarly simple yet a tighter lower bound alternative, is an extension to the multi-sample case (Burda et al., 2015). Indeed, consider the $n$-sample Jensen’s lower bound
$\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\coloneqq\mathbb{E}_{(c_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\!\left[\log\!\left(\!\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{i})\!\right)\!\right]\!.$ (6)
Importantly, the $\log$ function is outside of the $n$-sample average to tighten the bound. It is straightforward to verify that $\mathcal{L}_{\theta}^{(1)}(x,a^{\ast})$ recovers the Jensen’s lower bound as defined before in Eqn (4). As shown in Burda et al. (2015), the lower bound becomes tighter as $n$ increases $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\leq\mathcal{L}_{\theta}^{(n+1)}(x,a^{\ast})$ for any $n\geq 0$. As $n\rightarrow\infty$, the bound approaches the marginal likelihood $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\rightarrow\log\pi_{\theta}(a^{\ast}|x)$, the ultimate objective of interest, under certain regularity conditions on $\pi_{\theta}$.
To maximize the multi-sample lower bound $\mathcal{L}_{\theta}^{(n)}(x,a^{*})$ with gradient ascent, we can construct the REINFORCE stochastic gradient estimate as follows,
$\underbrace{\sum_{i=1}^{n}\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)\cdot\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)}_{g_{1}^{(n)}}$
$+\underbrace{\nabla_{\theta}\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{i})}_{g_{2}^{(n)}}.$ (7)
Empirically, the first term $g_{1}^{(n)}$ tends to have high variance as $n$ increases (Rainforth et al., 2018), since the objective $\log\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})$ correlates updates to all $n$ samples. Akin to before, we can introduce the leave-one-out control variate without incurring any bias for variance reduction (Mnih and Rezende, 2016; Kool et al., 2019)
$\sum_{i=1}^{n}\left(\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i}\right)\cdot\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$
where $\tilde{v}_{i}=\log\frac{1}{n-1}\sum_{j\neq i}\pi_{\theta}(a^{\ast}|x,c_{j})$.
Note that the second term $g_{2}^{(n)}$, though can be estimated via random samples, is unlike a regular SFT loss since it is the log average of multiple probabilities, instead of the average of log probabilities. As $n\rightarrow\infty$, since $\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{*}|x,c_{i})\rightarrow\log\pi_{\theta}(a^{*}|x)$, we see that at least conceptually $g_{2}^{(n)}$ can be understood as directly maximizing the marginal likelihood - the average over probabilities effectively marginalize the chain-of-thought conditional distribution.
As we will show in Section 8, multi-sample lower bound generally improves the single sample Jensen’s lower bound. This means that tightened lower bound improve training objectives both in theory and in practice.
1:  INPUT policy $\pi_{\theta}$
2:  while  $t=0,1,2...$  do
3:     (i) For each sampled prompt $x$, collect $n$ generations $(y_{i})_{i=1}^{n}$ and extract their corresponding chain-of-thoughts $(c_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)$.
4:     (ii) Evaluate $\pi_{\theta}(a^{*}|x,c_{i})$ with a forward pass; calculate gradients $\nabla_{\theta}\log\pi_{\theta}(c_{i}),\nabla_{\theta}\log\pi_{\theta}(a^{*}|x,c_{i})$ with backprop.
5:     (iii) Update $\theta$ with $n$-sample average of gradient estimate Eqn (5) or its multi-sample variant Eqn (7).
6:  end while
Algorithm 1 Chain-of-thought optimization with Jensen’s lower bound (or its multi-sample extension)
5 Connections to algorithmic alternatives
The lower bound approach bears close connections to a number of algorithmic alternatives, which we discuss below. See Algorithm 1 for the pseudocode for of full algorithm.
5.1 Evidence lower bound
The evidence lower bounds (ELBO) (Blei and Jordan, 2006; Kingma and Welling, 2013a; Burda et al., 2015) control for the tightness of the lower bound with an inference distribution $q_{\phi}(c|x,a^{\ast})$ which defines a distribution over chain-of-thoughts. The lower bound takes the following form
$\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathbb{E}_{c}\left[\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right],$
where the expectation is under $c\sim q_{\phi}(\cdot|x,a^{\ast})$. It lower bounds the marginal log likelihood $\mathcal{L}_{\theta,\phi}(x,a^{\ast})\leq\log\pi_{\theta}(a^{\ast}|x)$ and it is tight if and only if the inference distribution equals the posterior distribution $q_{\phi}(c|x,a^{\ast})=p^{\pi_{\theta}}(c|x,a^{\ast})$. Since ELBO is a function of both policy parameter $\theta$ and inference distribution parameter $\phi$, we can optimize both with stochastic gradient estimates: given a chain-of-thought sample $c\sim q_{\phi}(\cdot|x,a^{\ast})$,
$g_{\theta}$ $=\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x,c)+\nabla_{\theta}\log\pi_{\theta}(c|x),$
$g_{\phi}$ $=\nabla_{\phi}\log q_{\phi}(c|x,a^{\ast})\left(\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right)$
$\ \ -\nabla_{\phi}\log q_{\phi}(c|x,a^{*}).$
Juxtaposing the form of the gradient here and the gradient to the Jensen’s lower bound defined in Eqn (5), we observe that the inference distribution gradient $g_{\phi}$ bears resemblance to the REINFORCE gradient; while the policy distribution gradient $g_{\theta}$ bears resemblance to the SFT gradient. In fact, we can show that under the special parameterization $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, the two gradients are exactly equivalent. Such an observation is stated formally below.
Lemma 1.
(Jensen’s lower bound as a special case of ELBO) When $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, ELBO is equivalent to the Jensen’s lower bound $\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathcal{L}_{\theta}(x,a^{\ast})$ gradient esimtates are equivalent to the Jensen’s lower bound’s stochastic gradient estimates.
Proof.
The proof follows from the fact that when $q_{\phi}=\pi_{\theta}$, we have
$g_{\phi}=\nabla_{\theta}\log\pi_{\theta}(c|x)\cdot\log\pi_{\theta}(a^{\ast}|x,c)-\nabla_{\theta}\log\pi_{\theta}(c|x)$
Adding this gradient to $g_{\theta}$, a simple manipulation shows that the aggregate gradient is equivalent to the gradient of the lower bound defined in Eqn (5). ∎
By introducing an inference distribution $q_{\phi}$, ELBO is much expressive than the Jensen’s lower bound and allows for a tighter approximation to the marginal log likelihood. However, this also introduces additional complexity of having to learn the approximate posterior distribution. In our applications of interest, training a posterior model of a large size can be a major computational overhead. In practice, Hoffman et al. (2024) approximates the posterior via a few steps of MCMC and forgoes learning with such a distribution altogether. We take a different approach with a similar motivation: by tightening the lower bound with multiple samples, we also avoid the need for an explicit posterior.
5.2 Reinforcement learning
We show that there is a close connection between the lower bound formulation and the expected reward (return) maximization objective in RL (Sutton and Barto, 1998). Concretely, we will see how the lower bound objectives are closely related to a conditional expectation trick that produces a RL policy gradient estimate with lower variance. First, we show that (up to a log transform) RL optimizes for the same target as the lower bound objectives, given the indicator reward.
Lemma 2.
(RL optimality is equivalent to maximum likelihood optimality) When $r(x,y)=\mathds{1}_{\{a=a^{\ast}\}}$, the optimal policy to the RL objective is equivalent to the optimal policy of the maximum likelihood objective Eqn (3).
Proof.
The conclusion follows from the fact that $\mathbb{E}\left[\mathds{1}_{\{a=a^{\ast}\}}\right]=\pi_{\theta}(a^{*}|x)$. Hence the two objectives differ by a $\log$ operation and yield the same optimal solution. ∎
Assuming access to $n$ i.i.d. trajectories $(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)$, we start with the classic RL policy gradient with leave-one-out baseline (i.e., RLOO (Ahmadian et al., 2024b))
$g_{\text{vanilla-pg}}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\cdot\left(\mathds{1}_{\{a_{i}=a^{*}\}}-w_{i}\right),$ (8)
where $w_{i}=\frac{1}{n-1}\sum_{j\neq i}\mathds{1}_{\{a_{j}=a^{*}\}}$ is the leave-one-out baseline. Now, we present a new policy gradient estimate of the RL objective with guaranteed variance reduction, which is also feasible to implement with sample-based learning.
Definition 3 (A variance-reduced RL policy gradient estimate).
Given $n$ trajectories $(y_{i})_{i=1}^{n}$ from a single prompt $x$, we define $g_{\text{var-reduced-pg}}$ as
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot(\pi_{\theta}(a^{*}|c_{i})-\tilde{w}_{i})+\nabla_{\theta}\pi_{\theta}(a^{*}|c_{i}),$ (9)
where $\tilde{w}_{i}=\frac{1}{n-1}\sum_{j\neq i}\pi_{\theta}(a^{*}|c_{j})$ is the leave-one-out baseline akin to similar constructs in the lower bound case.
We show that the variance-reduced policy gradient estimate is closely related to the classic gradient estimate via the conditional expectation trick.
Lemma 4.
(Conditional expectation) Under the same assumption as Lemma 2 and denoting $a\sim\pi_{\theta}(\cdot|c)$ as the sampling process $a_{i}\sim\pi_{\theta}(\cdot|c_{i})$, it holds that $g_{\text{var-reduced-pg}}$ is a conditional expectation of $g_{\text{vanilla-pg}}$
$g_{\text{var-reduced-pg}}=\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[g_{\text{vanilla-pg}}\;|\;(c_{i})_{i=1}^{n}\right].$ (10)
We note that without the leave-one-out baselines $\tilde{w}_{i},\tilde{w}_{i}$, the conclusion Eqn (10) is straightforward as both estimates Eqn (9) and Eqn (8) become plain averages of i.i.d. terms. Now, using Lemma 4, we immediately see that the new gradient estimate yields smaller variance.
Theorem 5.
(Variance reduction) Under the same assumption as Lemma 2, we have guaranteed variance reduction
$\mathbb{V}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{var-reduced-pg}}\right]\leq\mathbb{V}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{vanilla-pg}}\right].$ (11)
The proof is provided in Appendix C. Putting $g_{\text{var-reduced-pg}}$ from Eqn (9) and the gradient estimate of the Jensen’s lower bound (Eqn (5)) side-by-side, we identify intriguing similarities. Both gradient estimates employ two terms that update either the chain-of-thought component $\pi_{\theta}(\cdot|x)$ or the answer component $\pi_{\theta}(\cdot|x,c)$, with the only subtle difference being the extra log-transform needed for obtaining the Jensen lower bound. This alludes to the fact that the lower bound gradient has intrinsic built-in variance reduction.
We provide additional discussions of a few practical trade-offs in using the variance-reduced estimate $g_{\text{var-reduced-pg}}$ in Appendix C.
Figure 2: Short-form answer experiments with MATH. We compare three baselines: online RL with access to the oracle Sympy-based reward and lower bound algorithms. In the left plot, we monitor the reward on the training dataset. Online RL obtains the best training time trade-off, followed by multi-sample lower bound and the single-sample lower bound; In the middle plot, we monitor the evaluation on a test set during training. Multi-sample lower bound and online RL obtains similar performance; In the right plot, we graph training reward against the lower bound objectives, averaged over training tokens. The two signals bear positive correlations overall and multi-sample lower bound yields better correlations.
5.3 Optimizing Jensen’s lower bound with regularization is optimizing a special ELBO
When optimzing the lower bound objectives, we also apply the KL regularization motivated from the regularized RL formulation (Eqn (1)). Though this combination seems ad-hoc, we will see that optimizing such an hybrid objective is in fact equivalent to maximizing a special ELBO.
Incorporating the regularization into the lower bound formulation, we have an aggregate objective
$\mathcal{L}_{\theta}(x,a^{*})-\beta\mathbb{KL}(\pi_{\theta},\pi_{\text{ref}}).$ (12)
If we refine the regularization a little more: instead of the generation level regularization, we apply regularization at the chain-of-thought: $\mathbb{KL}_{c}\left(\pi_{\theta},\pi_{\text{ref}}\right)\coloneqq\mathbb{E}_{c\sim\pi_{\theta}(\cdot|x)}\left[\log\frac{\pi_{\theta}(c|x)}{\pi_{\text{ref}}(c|x)}\right]$, then the resulting aggregate objective can be interpreted in a more coherent way, as an ELBO to a concrete generative process.
Lemma 6.
(Regularized lower bound as an ELBO to a special generative process) Assume a generative process $c\sim\pi_{\text{ref}}(\cdot|x),a\sim\pi_{\theta}(\cdot|x,c)$ that defines a marginal distribution $p_{\pi_{\theta},\pi_{\text{ref}}}(a|x)\coloneqq\sum_{c}\pi_{\text{ref}}(c|x)\pi_{\theta}(a^{*}|x,c)$. Then the regularized objective $\mathcal{L}_{\theta}(x,a^{*})-\mathbb{KL}_{c}(\pi_{\theta},\pi_{\text{ref}})$ is a lower bound to the log likelihood $\log p_{\pi_{\theta},\pi_{\text{ref}}}(a|x)$.
Proof.
Applying the same derivation as the regular ELBO, log likelihood $\log p_{\pi_{\theta},\pi_{\text{ref}}}(a|x)$ is lower bounded as
$\geq\max_{\phi}\ \mathbb{E}_{c\sim q_{\phi}(\cdot|x,a^{\ast})}\left[\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\text{ref}}(c|x)}\right]$
$\geq_{(a)}\mathbb{E}_{c\sim\pi_{\theta}(\cdot|x)}\left[\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{\pi_{\theta}(c|x)}{\pi_{\text{ref}}(c|x)}\right]$
$=\mathcal{L}_{\theta}(x,a^{*})-\mathbb{KL}_{c}(\pi_{\theta},\pi_{\text{ref}}),$
where inequality (a) is due to choosing $q_{\phi}=\pi_{\theta}$ and the last equality is by definition. Hence the proof is complete. ∎
Note that the aggregate objective Eqn (12) can also be optimized via stochastic gradient ascent. We just need to add an additional term associated with the KL regularization, to the original gradient estimate to $\mathcal{L}_{\theta}(x,a^{*})$ defined in Eqn (5). An example of such a gradient estimate usually takes the following form
$\log\frac{\pi_{\theta}(c|x)}{\pi_{\text{ref}}(c|x)}\nabla_{\theta}\log\pi_{\theta}(c|x),c\sim\pi_{\theta}(\cdot|x).$
Though our lower bound interpretation (Lemma 6) is under a regularization only on the chain-of-thought, in practice, we still apply the full generation level regularization following common practice (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022).
6 Lower bound vs. RL: practical trade-offs
In this section, we will discuss a few practical trade-offs of different algorithms.
6.1 Choice of reward
A key difference between RL and lower bound methods is the choice of reward. As discussed earlier, lower bound methods do not require a reward - more concretely, they can be understood as adopting the indicator reward $r(x,y)=\mathds{1}_{\{a=a^{\ast}\}}$. In practice, this can be instantiated as a strict string match float(answer == gt_answer). However, such a reward function will likely induce false negatives, as semantically equivalent generations might be vastly different strings. In practice, a more lenient match is typically applied to better balance the false negative. For example, for math problems (Hendrycks et al., 2021; Yue et al., 2024), usually programmtic checks are implemented to check for equivalence of two short expressions, such that e.g., pi and 3.1415926 might be considered equivalent.
More formally, consider a general reward function $r(x,y)=\text{match}(a,a^{*})$ calculated as a binary match between $a$ and $a^{*}$. We can rewrite the RL objective as $\mathbb{E}[\text{match}(a,a^{*})]$. In order to adapt the formulation in this work to the lower bound case, we need to work explicitly with the equivalent set $\mathcal{A}\coloneqq\{a|\text{match}(a,a^{*})=1\}$. We will need to calculate quantities such as the probability $\pi_{\theta}(\mathcal{A}|x,c)\coloneqq\sum_{a\in\mathcal{A}}\pi_{\theta}(a|x,c)$, which reduces to $\pi_{\theta}(a^{*}|x,c)$ in case we use exact match. Computing such probabilities is expensive since we need to enumerate all $a\in\mathcal{A}$ if inverting the match function is feasible at all. As a result, the lower bound formulation cannot be adapted to generic match function or reward function.
In summary, when a good reward is available, online RL is at an advantage due to its flexibility. There are also cases where good rewards are not easy to come by, and adopting the lower bound formulation is a decent default strategy. An example is where the ground truth answer takes a rather long form, in which case string match or programmatic check will produce too much false negatives.
6.2 Further details of lower bound algorithms
One technical challenge of the lower bound algorithm is that, given a full generation $y$, we need to define the boundary between the chain-of-thought $c$ and answer $a$ such that $y=[c,a]$ where we can interpret the latter as the concatenation of two arrays of tokens (or strings). Usually, we can ask the model to generate responses in a particular format, e.g., concluding the generation with a phrase the final answer is (Yue et al., 2024). We can use this phrase as an identifier, and define everything after the phrase as the answer $a$ and otherwise as the chain-of-thought (this step is part of the algorithm pseudocode in Algorithm 1). During training, we will need to enforce such formatting from the model so that the loss can be applied properly.
One mitigating strategy is to apply the KL regularization against the reference policy. The initial motive for such regularization was to prevent the model from degrading its general capability during RL finetuning (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). We will ablate how useful such regularization is to the lower bound methods. Besides regularization, we can also add a formatting penalty to the model that encourages following the format. Echoing the previous discussion on the reward choice, we highlight here that in general checking if the model complies with the formatting requirement is much simpler than checking if the answers are matched.
7 Related work
Chain-of-thought as latent variable modeling.
The idea of casting optimizing chain-of-thought as latent variable modeling is not new. Previously, Hoffman et al. (2024) proposed an algorithm motivated by maximizing ELBO to tackle reasoning problems. Such an algorithm also draws close connections to prior work (Zelikman et al., 2022; Gulcehre et al., 2023; Singh et al., 2023; Yuan et al., 2023) all of which resemble a hybrid offline-online RL training loop, where they alternate between sampling and filtering via a reward. They also have an interpretation as EM algorithmic variants (Moon, 1996).
Despite the appeal of a full ELBO formulation, it is rarely implemented in practice due to the requirement of learning the posterior distribution. Indeed, despite the formulation of Hoffman et al. (2024) they ended up approximating the posterior with MCMC, which effectively made use of an explicit reward to filter samples. This also marks a key difference from our work - we do not apply any explicit reward scoring throughout our algorithmic design and practical implementation. In addition, Hu et al. (2024) has proposed a more systemic hierarchical latent variable modeling view of chain-of-thought. Similar to our motive, Sordoni et al. (2024) optimized an ELBO inspired objective for prompt selection, where they did not resort to an external reward.
Evidence lower bound and RL.
The connections between evidence lower bound and RL has been extensively studied in both the variational inference (Ranganath et al., 2016; Blei et al., 2017) and RL community (Levine, 2018; O’Donoghue et al., 2020). In the RL literature, much of the variational inference view has often been harnessed to better interpret existing algorithms with much focus on the goal-conditional problems, where a single reward is assigned at the end of a trajectory, akin to the RLHF case (Andrychowicz et al., 2017; Eysenbach et al., 2020; Tang and Kucukelbir, 2021). Our formulation also naturally incorporates the tighter multi-sample lower bound (Burda et al., 2015; Rainforth et al., 2018) as special cases, which has seen little adoption in prior RL literature.
Figure 3: Ablation of number of samples $n$ for multi-sample lower bounds. As we increase the number of samples, the multi-sample lower bound seems to further improve the training-time efficiency. This corroborates the theoretical insight that as $n$ increases, the multi-sample lower bound objectives become tighter.
8 Experiments
We empirically validate the relative strength of the lower bound methods against online RL algorithm. We focus on the mathematical reasoning dataset MATH (Hendrycks et al., 2021) where the prompt $x$ asks the model a mathematical question with a short form answer $a^{*}$. We mainly focus on the two algorithmic variants proposed in this work: the lower bound algorithm defined through the gradient estimate in Eqn (5) as well as its multi-sample variant Eqn (7). As a strong baseline, we consider the online policy gradient RL algorithm which applies Sympy (Meurer et al., 2017) to automatically match the answers. The RL algorithm applies leave-one-out for variance reduction, as is commonly practiced (Ahmadian et al., 2024a; Shao et al., 2024). Our main experiments are based on the 8B and 70B model from the Llama 3 model family (Dubey et al., 2024). All algorithmic variants apply identical hyper-parameters such as learning rate, and that they all apply $n=4$ samples for gradient estimations, which we detail in Appendix B.
We highlight again that the RL baseline is at an advantage in this setting, since the reward is fairly accurate and is itself being used as evaluation signals too (Yue et al., 2024). We do not compare with other baselines developed in prior work (e.g., (Hoffman et al., 2024)) as they can be interpreted as variants of online RL algorithms with certain low-level implementation differences.
8.1 Comparion on MATH
During RL training, we use a reward of $r=1$ when there is an answer match and $r=0$ otherwise. Note that lower bound algorithms does not require access to such a reward, but we monitor the reward scores during training. Figure 2 left plot shows the training performance of all baselines. For the x-axis, we use the KL divergence $\mathbb{KL}(\pi_{\theta},\pi_{\text{ref}})$ calculated over the training set. Following the practice in (Gao et al., 2023), we adopt the KL divergence as a certain measure of the optimization budget that the algorithm has consumed. Note that here all experiments are run with the same regularization coefficient $\beta=10^{-3}$ since it achieves a good trade-off for all algorithmic variants over all.
Training performance.
Figure 2 left plot shows that online RL achieves a good KL-performance trade-off on the training set. This is probably not a big surprise since online RL optimizes for the very same objective that we monitor here. In the meantime, lower bound algorithms enjoy reasonable performance: as the policy deviates from the reference policy, the reward performance improves despite not explicitly training for it (in theory lower bound algorithms optimize for a hard string match rather than Sympy match). (2) the multi-sample lower bound algorithm obtains noticeably better performance than the one-sample lower bound baseline, despite using the same $n=4$ generations per update. We will ablate on the impact of parameter $n$ on the performance.
Evaluation.
Figure 2 middle plot shows the evaluation performance on an held-out test set. We note that the reward on the training set is higher than the test set, because the model has been SFT’ed on on the training prompts. For evaluation, observe that the multi-sample lower bound method obtains similar performance as online RL, despite being outperformed during training. We conjecture that this is because online RL tends to overfit the training prompts more significantly, producing a high training reward that does not transfer as well to the evaluation time. This shows that even without training on the reward signal explicitly, lower bound algorithms can obtain a similar evaluation performance as online RL.
Statistical correlation between objectives.
Figure 2 right plot graphs the training time reward against the lower bound objectives. If we consider the training reward as a ground truth objective to optimize for, we see that the multi-sample lower bound displays a stronger correlations between the surrogate objective and the ground truth. This corroborates with the observation that multi-sample lower bound tends to lead to better performance, compared to single-sample lower bound.
8.2 Ablation study
We now provide ablation results on a few important dimensions of the algorithm.
Figure 4: Ablation of regularization coefficient $\beta$. As $\beta$ increases, all algorithmic variants seem to obtain better efficiency in the training performance-KL divergence trade-off. However, strong regularization also prevents the policy from deviating much from the reference policy, preventing bigger training improvements.
Multi-sample ablation on sample size $n$.
We ablate on the number of sample $n$ used for constructing per gradient update. In theory, as $n$ increases, the multi-sample lower bound becomes tighter and asymptotically approaches the marginal likelihood objective (which is equivalent to the RL objective). We vary the sample size $n\in\{1,2,4,8\}$ and compare the performance. Figure 3 shows that as $n$ increases, the algorithm becomes more KL-efficient: with a fixed budget on KL, the model obtains better performance. Intriguingly, we also observe a training performance akin to reward over-optimization (Gao et al., 2023) - as the optimization progresses, the training reward drops slightly (for blue curve). We can interpret this as the result of the fact that lower bound algorithms do not optimize for the same indicator matching function as the reward we monitor.
Regularization ablation.
We investigate the impact of the regularization coefficient $\beta\in\{0,10^{-3},10^{-2},10^{-1}\}$. Figure 4 shows the training performance of the single-sample lower bound vs. online RL. One observation is that as $\beta$ increases, the trade-off efficiency for both algorithms improves - however, in general the algorithm also makes less deviation from the reference policy, hence leading to less improvement for a fixed training steps.
Figure 5: Ablation of model size (8B vs. 70B). We find that the multi-sample lower bound algorithm is fairly competitive against the online RL algorithm in the 70B scale. Both algorithm traces out a similar KL-performance trade-off, with multi-sample lower bound algorithm obtaining a slightly better performance given a similar compute budget as online RL.
Scaling up model size.
Since the multi-sample lower bound algorithm appears more competitive, we compare it against the online RL in the 70B case. Figure 5 shows that the lower bound algorithm obtains competitive performance against online RL in terms of the KL-performance trade-off. With roughly the same amount of compute budget, we find that the lower bound algorithm seems to drift further from the reference policy, hence extending the trade-off curve to a performance of $70\%$ test set accuracy, which outperforms online RL modestly.
9 Conclusion and limitations
While the theoretical perspective that views chain-of-thought as latent variable is not new, we propose a much practically simpler approach to maximizing the evidence lower bound. The lower bound technique draws intriguing connections between SFT, RL, and sheds light on how literature on probabilistic inference produces algorithmic insights to optimizing chain-of-thought.
Possible directions for future research include studying how such algorithms can be applied to domains where good reward is difficult to come by, and how they scale with data quality compared to RL, how the approach might generalize to hidden chain-of-thought (Jaech et al., 2024; Xiang et al., 2025), and how the algorithm might combine with a reward function or verifier that generates ground truth answer in an iterative fashion.
APPENDICES: Learning to chain-of-thought with Jensen’s evidence lower bound
Appendix A A review of the graphical model perspective
We make a more extended discussion about the graphical model shown in Figure 1.
Probabilistic inference with a learnable prior.
Figure 1(a) shows the generic structure for probabilistic inference with a learnable prior, with latent variable $z$ and observable $o$. Here, $\theta$ controls both the prior and observation generation:
$z\sim p_{\theta}(\cdot),o\sim p_{\theta}(\cdot|c).$
The inference parameter $\phi$ denotes a the posterior inference distribution $q_{\phi}(z|o)$ that seeks to approximate the true posterior $p_{\theta}(z|o)\coloneqq\frac{p_{\theta}(c)p_{\theta}(o|c)}{\sum_{c^{\prime}}p_{\theta}(c^{\prime})p_{\theta}(o|c^{\prime})}$. Together, they can form an ELBO that lower bounds the marginal log likelihood [Blei et al., 2017]
$\log p_{\theta}(o)\geq\underbrace{\mathbb{E}_{z\sim q_{\theta}(\cdot|o)}\left[\log p_{\theta}(z|o)+\log\frac{q_{\phi}(z|o)}{p_{\theta}(z)}\right]}_{\mathcal{L}_{\theta,\phi}(o)}.$
The right hand side $\mathcal{L}_{\theta,\phi}(o)$ can be optimized via stochastic gradient descent on the joint variable $(\theta,\phi)$. The lower bound is tight when the inference distribution is exactly the posterior $q_{\phi}(z|o)=p_{\theta}(z|o)$. A learnable prior refers to the fact that the prior distribution over latent $p_{\theta}(z)$ depends on $\theta$ too, while in much of the prior literature is is kept constant [Hoffman and Johnson, 2016, Blei et al., 2017]. For the transition from generic probabilistic inference to our use case, a learnable prior is also fundamentally important.
Chain-of-thought with full ELBO.
Figure 1(b) shows a direct mapping of the probabilistic inference structure to the case of optimizing chain-of-thought. Here, the chain-of-thought $c$ is the latent variable and the ground truth answer $a^{*}$ is the observable. A more precise mathematically definition would be to consider yet another binary optimality variable $O=\mathds{1}_{\{a=a^{\ast}\}}$ that determines whether the random variable answer $a$ is optimal. Here, we directly replace it with $a^{*}$ for notational simplicity.
If we further introduce a general conditional dependency on the prompt $x$, we arrive at the lower bound defined in Eqn (4)
$\log\pi_{\theta}(a^{*}|x)\geq\underbrace{\mathbb{E}_{c\sim q_{\theta}(\cdot|x,a^{*})}\left[\log\pi_{\theta}(c|x,a^{*})+\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right]}_{\mathcal{L}_{\theta,\phi}(x,a^{*})}.$
Chain-of-thought with Jensen’s lower bound.
In Figure 1(c), we replace the variational posterior $q_{\phi}$ by the prior distribution itself $\pi_{\theta}$. As discussed in the main paper, this looses the lower bound but make the optimization objective much simpler. See detailed derivations in Section 3. We see there there appears to be a duplicated arrow that goes from $\theta$ to the latent variable $c$. We make such duplication to distinguish between the inference distribution (dashed arrow) and the generative distribution (solid arrow); in this particular case, we deliberately make the two distributions identical.
Jensen’s lower bound with regularization.
Finally, Figure 1(d) presents the graphical model for the case where a KL regularization is added to the Jensen’s lower bound (see Lemma 6 for formal statements). In this case, the generative prior distribution is computed from the reference policy $\pi_{\text{ref}}$ parameterized by $\theta_{\text{ref}}$ which is kept fixed during training, while the rest of the distributions are still parameterized by $\theta$.
Appendix B Hyper-parameters and experimental settings
[…]
We train on the MATH training set with $7500$ examples and evaluate on the test set with $2500$ examples. A supervised fine-tuning on the training set is conducted to warm up the RL training, hence the gap between training and test set accuracy.
For both training and evaluation, we provide system instructions that ask the model to generate a response with step-by-step solution, followed by a final conclusion phrased as the final answer is followed by the answer predicted by the model. This is consistent with the prompt structure discussed for Llama models [Dubey et al., 2024, Yue et al., 2024].
All experiments are conducted with an entropy regularization coefficient $\beta>0$ which we have ablated in the main paper.
B.1 Lower bound algorithm details
We add an additional RL loss with the reward as $r_{\text{reg}}(x,y)=-10$ if $y$ does not follow the formatting requirement (that the identifier phrase the final answer is is in $y$) and zero otherwise. We find that this generally helps stabilize the lower bound training process. This is especially useful for small models where under temperature sampling, it can often not follow instructions strictly.
The lower bound loss is applied only to generations that abide by the formatting requirement, i.e., with the final answer is found inside the full generation. Otherwise, the loss is masked out. After generating a response $y$ from the model, we can parse out its chain-of-thought component $y=(c,a)$ and replace its generated answer with the ground truth. This produces another sequence $y^{*}=(c,a^{*})$. Running a forward pass on such a sequence produces quantities such as $\log\pi_{\theta}(c|x)$ and $\log\pi_{\theta}(a^{*}|c,x)$ required for the algorithm.
Advantage post-processing.
Both the baseline RL and the lower bound algorithm applies advantage post-processing, following common practice [Schulman et al., 2017]. For example, in the multi-sample lower bound algorithm, the advantage for the $i$-th generation is
$A_{i}=\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i}.$
A further normalization is applied to the advantage $\tilde{A}_{i}=\text{clip}(A_{i}/\text{std}\left(A\right),-1,1)$ such that the outcome $\tilde{A}_{i}$ is applied in the actual update. For the baseline RL algorithm, a similar procedure applies. Put together, the fully implemented multi-sample lower bound update is
$\frac{1}{n}\sum_{i=1}^{n}\left(\left(\tilde{A}_{i}+\tilde{A}_{i}^{\text{(ref)}}\right)\nabla_{\theta}\log\pi_{\theta}(a_{i}|x,c_{i})\right)+\nabla_{\theta}\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{i})-\beta\nabla_{\theta}\mathbb{KL}\left(\pi_{\theta}(\cdot|x),\pi_{\text{ref}}(\cdot|x)\right),$
where $\tilde{A}_{i}^{\text{ref}}=\text{clip}\left(\left(r_{\text{ref}}(x,y_{i})-\bar{r}_{\text{ref}}\right)/\text{std}(r_{\text{reg}}),-1,1\right)$ is the normalized advantage for the formatting penalty. The normalization makes it such that the ultimate update optimizes for a lower bound more akin to the weighted lower bound [Higgins et al., 2017] though the underlying algorithmic motivations differ.
Appendix C Variance-reduced RL gradient estimate
We provide more discussion on the variance-reduced RL gradient estimate.
C.1 Proof of variance reduction
Recall that we denote $(y_{i})_{i=1}^{n}$ as the set of generations and $(c_{i})_{i=1}^{n}$ be the set of chain-of-thoughts generated from prompt $x$. We drop the dependency on prompt $x$ wherever the context is clear.
Proof of Theorem 11.
A direct computation shows that
$\begin{split}\mathbb{V}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{vanilla-pg}}\right]=&\mathbb{E}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{vanilla-pg}}-g_{\text{var-reduced-pg}}+g_{\text{var-reduced-pg}}-\mathbb{E}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{vanilla-pg}}\right]\right]\\ =&\mathbb{E}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[\|g_{\text{vanilla-pg}}-g_{\text{var-reduced-pg}}\|^{2}\right]+\mathbb{V}_{(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)}\left[g_{\text{var-reduced-pg}}\right],\end{split}$ (13)
where the cross-term vanishes due to Eqn (10). From this, Eqn (11) follows immediately. ∎
Proof of Lemma 4.
We begin by computing the conditional expectation $\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[g_{\text{vanilla-pg}}\;|\;(c_{i})_{i=1}^{n}\right]$, which yields
$\begin{split}\underbrace{\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(y_{i})\cdot\mathds{1}_{\{a_{i}=a^{*}\}}\;|\;(c_{i})_{i=1}^{n}\right]}_{\text{I}}+\underbrace{\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(y_{i})\cdot\tilde{w}_{i}\right]}_{\text{II}}.\end{split}$ (14)
where we use the notation $a\sim\pi_{\theta}(\cdot|c)$ to indicate that each answer $a_{i}\sim\pi_{\theta}(\cdot|c_{i})$ is i.i.d. sampled from its corresponding chain-of-thought. Expanding the first term I, we have
$\begin{split}\text{I}=_{(a)}&\frac{1}{n}\sum_{i=1}^{n}\sum_{a}\big{(}\nabla_{\theta}\log\pi_{\theta}(a|c_{i})+\nabla_{\theta}\log\pi_{\theta}(c_{i})\big{)}\cdot\mathds{1}_{\{a=a^{*}\}}\cdot\pi_{\theta}(a|c_{i})\\ =_{(b)}&\frac{1}{n}\sum_{i=1}^{n}\big{(}\nabla_{\theta}\pi_{\theta}(a^{*}|c_{i})+\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot\pi_{\theta}(a^{*}|c_{i})\big{)},\end{split}$ (15)
where (a) is by definition of the expectation and $a\in\mathcal{A}$ denotes a dummy answer variable; (b) is due to the definition of the indcator function. Now recalling the definition of $w_{i}$ as leave-one-out baseline to simplify term II:
$\begin{split}\text{II}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[\nabla_{\theta}\log\pi_{\theta}(y_{i})\cdot w_{i}\;|\;(c_{i})_{i=1}^{n}\right]=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}\left[\nabla_{\theta}\log\pi_{\theta}(y_{i})\cdot\mathds{1}_{\{a_{j}=a^{*}\}}\;|\;(c_{i})_{i=1}^{n}\right].\end{split}$ (16)
Note we can explicitly compute each summand on the right hand side of Eqn (16) as product of two conditional expectations, thanks to the fact that when $i\neq j$:
$\begin{split}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|c)}[\nabla_{\theta}\log\pi_{\theta}(y_{i})\cdot\mathds{1}_{\{a_{j}=a^{*}\}}\;|\;(c_{i})_{i=1}^{n}]&=_{(a)}\big{(}\mathbb{E}_{a_{i}\sim\pi_{\theta}(\cdot|c_{i})}[\nabla_{\theta}\log\pi_{\theta}(a_{i}|c_{i})|c_{i}]+\nabla_{\theta}\log\pi_{\theta}(c_{i})\big{)}\cdot\pi_{\theta}(a^{*}|c_{j})\\ &=_{(b)}\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot\pi_{\theta}(a^{*}|c_{j}),\end{split}$ (17)
where (a) is due to the definition of the indicator function; (b) is based on the zero-mean property of score functions. Plugging Eqn (17) into the right hand side of Eqn (16), we have
$\begin{split}\text{II}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot\frac{1}{n-1}\sum_{j\neq i}\pi_{\theta}(a^{*}|c_{j})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot\tilde{w}_{i},\end{split}$ (18)
where we used the definition of $\tilde{w}_{i}$ from Eqn (9). Lastly, we combine Eqn (15) and Eqn (18) and obtain
$\begin{split}\text{I}+\text{II}=\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\nabla_{\theta}\pi_{\theta}(a^{*}|c_{i})+\nabla_{\theta}\log\pi_{\theta}(c_{i})\cdot\big{(}\pi_{\theta}(a^{*}|c_{i})-\tilde{w}_{i}\big{)}\bigg{)}=g_{\text{var-reduced-pg}}.\end{split}$ (19)
Thus we have concluded the proof of Lemma 4. ∎
C.2 Additional discussion on the practical trade-offs of the estimate
The form of the variance reduced gradient estimate implies that for the $i$-th generation, we have an advantage function of the form $\pi_{\theta}(a^{*}|c_{i})$. Though mathematically sound, since $a^{*}$ is made up of multiple tokens, the quantity $\pi_{\theta}(a^{*}|c)$ becomes exponentially small as the length of $a^{*}$ increases. As a result, we might need additional normalization technique to make the effective advantage function more suitable for numerical optimization.